Kids’ Candy Corporation (KCC), a large nationwide candy manufacturer, is considering the purchase of Holiday Chocolates, Inc. KCC’s management foresees much synergy between the two firms as a key outcome of a possible merger. As the Acquisitions Analyst employed by KCC, you have developed the following estimates for Holiday Chocolate’s incremental “Free Cash Flows” (FCF) for the upcoming four years:

Year

Free Cash Flow

1

$3.5 million

2

$5.9 million

3

$7.4 million

4

$8.8 million

              Your analysis further predicts that “Free Cash Flows” occurring after Year 4 will grow at a constant four percent (4%) annual rate. Holiday Chocolates, Inc. is estimated to have a “Weighted Average Cost of Capital” (WACC) of 8%, based on the capital structure that will be put in place following the acquisition, if it is made. During your employment at KCC, you have learned that the management team prefers the “Corporate Valuation Model” (CVM) method when evaluating possible acquisition candidates.

              You are preparing a presentation on this proposal for the management meeting tomorrow; the possible acquisition of Holiday Chocolates, Inc. is considered to be the most important agenda item for this meeting. In preparation for this meeting, you prepare answers to the following questions:

  1. Calculate the “Horizon Value” (HV) as of the end of Year Four for Holiday Chocolate’s expected “Free Cash Flows” that will occur after that date.
  1. What is the present value of the expected “Free Cash Flows” that Holiday Chocolates, Inc. is projected to generate, including the “Horizon Value,” given the assumptions outlined above?
  1. If KCC acquires only the assets and operations of Holiday Chocolates, In, what is the maximum offer that KCC should make when negotiating the purchase of Holiday Chocolates?

a. Why should KCC not pay more than this amount for Holiday Chocolates, Inc.?

  1. Now assume that Holiday Chocolates, Inc. has $12 million of debt outstanding. If KCC were to acquire the stock of Holiday Chocolates (i.e., an equity acquisition), what would then be the maximum amount that KCC should pay for Holiday Chocolates (i.e., what is Holiday Chocolates, Inc.’s “Net Asset Value?”)?
Answer #1
[$ in millions] 0 1 2 3 4
FCF $           3.50 $          5.90 $          7.40 $         8.80
PVIF at 8% 0.92593 0.85734 0.79383 0.73503
PV at 8% $           3.24 $          5.06 $          5.87 $         6.47
Sum of PV of FCF years 1 to 4 $        20.64
a] HV at the end of year 4 = 8.80*1.04/(0.08-0.04) = $     228.80
PV of HV = 228.80*0.73503 = $      168.17
b] PV of FCF [including PV of HV] $      188.82
c] Maximum payable for assets and operations = $      188.82
d] Maximum payable for stock = 188.82-12 = $      176.82